OpenStudy (anonymous):

-5x^3+7x^2-3x-2 x=-1

OpenStudy (anonymous):

x=-15

OpenStudy (anonymous):

after i do all the substitutions and steps im lost

OpenStudy (anonymous):

well -125+49-3x-2x=-1 then 76-5x = -1 -5x = 75 divide and get x =-15

OpenStudy (anonymous):

oops 76 should be negative up there

OpenStudy (anonymous):

-5(-1)^3+7(-1)^2-3(-1)-2 thats how i started out with mine.

OpenStudy (anonymous):

ohhh did you mean x in the equation = -1

OpenStudy (anonymous):

yea x mean -1

OpenStudy (anonymous):

after this step im lost. i know i use order of operations but the example they give me i dont get either lol

jhonyy9 (jhonyy9):

-5(-1)+7(1)-3(-1)-2=5+7+3-2=13